Circular No. 6680

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GK PERSEI S. Balman, H. B. Ogelman, and M. Orio, University of Wisconsin

at Madison, report: “In a pointed observation with the High Resolution Imager onboard the satellite ROSAT, the nova shell of the classical nova Per 1901 (GK Per) has been detected in the x-ray energy band 0.1–2.4 keV with an effective exposure of 53646.4 s. This is the first detection of a nova shell in x-rays. The x-ray nebula extends 46′′ southwest, 60′′ northwest, 52′′ southeast, and 43′′ northeast of GK Per, with an elliptical shape. The shell shows a clumpy morphology with several knots that are centrally bright and limb-darkened. The size of the knots vary between 4′′ and 8′′. The xray knots are not coincident with the optical knots; however, there is x-ray emission coincident with the radio maxima and a ridge in the southwestern part of the nebula. The surface brightness of the nebula is 703.4 ± 54.5 counts arcmin^{−2} (0.01 ± 0.001 counts/s), and the HRI countrate ratio for the point source and the shell indicates that only 20 percent of the x-ray flux is coming from the shell. ROSAT PSPC data show that the spectrum of the shell is thermal in origin.”

S/1997 (3671) 1 S. Mottola and G. Hahn, Institute of Planetary Exploration, German

Aerospace Research, Berlin; and P. Pravec and L. Šarounová, Ondřejov Observatory, report: “Our photometric observations of this Apollo object from the European Southern Observatory (La Silla) and Ondřejov on nine nights during May 30.9-June 8.1 UT revealed the presence of attenuations of the object’s brightness of ∼ 0.08 mag lasting for ∼ 2 hr, superimposed on its normal rotational lightcurve (P = 2.705 hr, amplitude 0.14 mag). The attenuation events were detected on May 30.985, June 1.131, 6.918, and 8.071; a period of 1.155 days was derived from the observations. We interpret these features as occultations/eclipses of a satellite orbiting (3671) Dionysus. The observed lightcurve characteristics are similar to those of 1994 AW_{1} (Pravec and Hahn 1997, Icarus 127, 431) and 1991 VH (IAUC 6607), suggesting that a binary nature may be relatively common among certain classes of near-earth asteroids. Further observations of the occultation/eclipse events are needed to model the dynamics of this system, and interested observers are encouraged to participate in a monitoring campaign. For updated information, see http://earn.dlr.de/dionysus.”

1997 June 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green